OpenStudy (anonymous):

Ok so I cant figure out why f(x) changes to 104 in this equation. f(x)=√x+17-5. Why?

OpenStudy (anonymous):

what do u mean f(x) changes to 104? is 104 the answer?

OpenStudy (anonymous):

and it's Sqrt x+17 but the -5 isn't part of the square root right?

OpenStudy (anonymous):

No ok so my math lab worked it out for me and this is how it did it. f(x)=√x+17-5 f(104)=√104+17-5 =√121-5 =11-6 =5. I wanna know why x changes to 104 I dont understand that.

OpenStudy (anonymous):

It is.

OpenStudy (anonymous):

ok, I get what u are saying, then the explanation is really simple. We are replacing x with 104.

OpenStudy (anonymous):

Yes but why? Lets say it where f(x)=√x+10-5 what wud x be now?

OpenStudy (anonymous):

Lets use an example, say f(x) = x + 3 and I want to figure out f(3) f(x) = x + 3 (3) = 3 + 3 = 6

OpenStudy (anonymous):

We are simply substituting or replacing the value for x.

OpenStudy (anonymous):

YOu can imagine f(x) is just y.

OpenStudy (anonymous):

I forgot the f on the second line. f(3) = 3 + 3 = 6***

OpenStudy (anonymous):

f(x)=√x+10-5 putting it into words, we want to replace x with x, so the equation will be the same. f(x)=√x+10-5 but if it is like this, f(3)=√x+10-5 it will become this. f(3)=√3+10-5 then solve.

OpenStudy (anonymous):

x is the independent variable and f(x) (or y) is dependent on it, so as JayDs said sub 104 for x and you will get y

OpenStudy (anonymous):

Man this isnt making sense.

OpenStudy (anonymous):

sorry guys.

OpenStudy (anonymous):

answer wold be 8

OpenStudy (anonymous):

do you not understand my clear examples?

OpenStudy (anonymous):

would I have to find the square root of 13?

OpenStudy (anonymous):

and subtract that from 8

OpenStudy (anonymous):

It is a simple substitution method. Maybe you should google it. f(x) is the same thing as y so saying f(x) = 3x + 5 is the same as y=3x + 5 f(x) = 3x + 5 f(3)- this means we want to replace x with 3. f(x) = 3(3) + 5 = 9 + 5 =14

OpenStudy (anonymous):

I would say unless specified saying "Round off to two decimal places or something else" leave it in exact form.

OpenStudy (anonymous):

sorry its cause on my mathlab it teaches me to do it another way

OpenStudy (anonymous):

is their more that 1 way to do it?

OpenStudy (anonymous):

*Than

OpenStudy (anonymous):

ok no problem but then their method must be really weird.

OpenStudy (anonymous):

no there isn't, it is simply a "Substitution method"

OpenStudy (anonymous):

f(x)=√x+17-5 f(104)=√104+17-5 =√121-5 =11-6 =5.

OpenStudy (anonymous):

hence the word "Substitute" if you know what that means.

OpenStudy (anonymous):

thats how they did it

OpenStudy (anonymous):

are u familiar with that method

OpenStudy (anonymous):

yes, it is the same thing as what I did.

OpenStudy (anonymous):

but ur answer is wrong, it is not 5 by 6.

OpenStudy (anonymous):

type it in a calculator and u will see that it is 6.

OpenStudy (anonymous):

he/she/they (your maths lab) or whatever u call it has made a big mistake.

OpenStudy (anonymous):

well for the last calculation part.

OpenStudy (anonymous):

yeah it was 6 my mistake

OpenStudy (anonymous):

I hope you understand the meaning and concept of "Substituting" it's basically saying I want whatever inside the bracket to be replaced with something else. f(x) = 3x <<say this is given. f(1) = ? we want to find this putting it in words, I want to replace "x" with 1 f(1) = 3(1) = 3

OpenStudy (anonymous):

Ok understood

OpenStudy (anonymous):

u know what happend i thought they werent giving me the value for x turns out they were lol. It was me that messed up.

OpenStudy (anonymous):

lol, u shouldn't say that.

OpenStudy (anonymous):

The most important thing is if you are stuck, ask ur math lab, teacher or whoever for classification.

OpenStudy (anonymous):

Ok thank you. You were very helpful

OpenStudy (anonymous):

no problem, as long as I could help u with ur problem.

OpenStudy (anonymous):

excuse the repetition of the word problem lol.